OpenStudy (anonymous):

HELP PLEASE Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). Answer y = 4/3x + 8 y = 4/3x + 3 y = -3/4 x + 8 y = - 3/4x + 3

OpenStudy (anonymous):

ok that makes more sense.

OpenStudy (anonymous):

slope is \[\frac{4}{3}\] then \[y-4=\frac{4}{3}(x+3)\] \[y-4=\frac{4}{3}x+4\] \[y=\frac{4}{3}x+8\]